# Adapting Large Language Models for Education: Foundational Capabilities, Potentials, and Challenges

### v2

@@ 2.4.3 Hybrid Methodologies for Reasoning
Table 1: Overview of LLMs’ performance on foundational education-related capabilities.
@@
GPT-4 (OpenAI, 2023b) 92.00 62.00 67.00 91.40 59.00 68.70
ChatGPT (OpenAI, 2023a) 57.10 48.60 48.10 79.50 47.00 54.40
TigerBot-70B-Chat-V2(70B) (TigerResearch, 2023) 54.36 61.30 30.50 82.83 75.40 -
LLaMA2(70B) (Touvron et al., 2023b) 60.27 51.60 29.90 82.30 56.18 55.20
LLaMA(65B) (Touvron et al., 2023a) 43.37 47.10 23.70 82.30 55.09 38.80
Yi(34B) (lingyiwanwu, 2023) 50.64 48.90 26.20 82.00 56.23 81.40
Vicuna(33B) (TheVicunaTeam, 2023) 13.72 44.90 15.20 83.00 56.16 39.80
WizardLM(30B) (Xu et al., 2023) 34.42 - 26.08 76.32 49.14 -
Moss(16B) (FudanUniversity, 2023) 6.90 39.00 - 55.80 49.00 33.10
Qwen(14B) (Bai et al., 2023) 58.98 52.7 43.90 80.20 49.43 72.10
Baichuan2(13B) (Yang et al., 2023a) 55.30 51.50 17.07 66.90 48.98 40.00
Alpaca(7B) (Taori et al., 2023) 0.15 39.50 9.10 75.71 36.28 29.90
ChatGLM3(6B) (Zeng et al., 2022) 72.30 43.10 44.50 76.50 - 69.00
2.5 Knowledge-based Question Answering
In the context of Knowledge-based Question Answering using LLM, the user presents a question to LLM, and LLM leverages knowledge-based methods and responds with the corresponding answer. Previous work by Ren et al. (2023) showed that LLMs have an inaccurate perception of factual boundaries and often exhibit overconfidence. Many studies have explored and utilized external knowledge from open-world and domain-specific databases to enhance the knowledge base of these LLMs.
2.5.1 Open-domain QA
Open-domain question answering requires LLMs to accurately determine the reliability of information in the open world and craft their responses based on that understanding. Jiang et al. (2021) evaluated the accuracy of LLM responses to a particular question from the perspective of calibration. Through experiments, the researchers discovered that models such as T5, BERT, and GPT-2 are not well-calibrated in QA tasks. While suggesting that incorporating calibration-related methods into the fine-tuning process can effectively enhance performance in QA tasks, it is evident that solely pre-trained language models still face significant challenges in open-domain tasks. To overcome this challenge, many works tried to add additional information to help the LLMs answer correctly (Khandelwal et al., 2019; Guu et al., 2020; Borgeaud et al., 2022). Lazaridou et al. (2022) employed information gathered from web searches as prompt input for LLM, conditioning it to generate answers to questions. This approach effectively enables LLM to use open-world information to answer questions.
2.5.2 Domain-specific QA
Although LLMs are trained on vast corpora, they may still exhibit gaps in understanding specific domains. Typically, there are dedicated knowledge repositories for professional content that consolidate domain-specific knowledge, such as MedlinePlus [^2] https://medlineplus.gov, GeeksforGeeks [^3] https://www.geeksforgeeks.org, etc. Choi et al. (2023) utilized an external knowledge base to generate a set of question-answer pairs and then employed fine-tuning to transfer financial knowledge to LLMs, significantly improving financial question-answering tasks. Another common approach is to leverage the in-context learning capability of LLMs by incorporating retrieved knowledge from the knowledge base into prompts. Peng et al. (2023a) used text embeddings for retrieval, followed by LLM’s automatic feature extraction in order to achieve higher accuracy in pest identification tasks. zhang et al. (2023d) utilized K-nearest neighbors (KNN) to search for the most similar K records from an accounting database, serving as k-shot examples and greatly improving accounting efficiency. There are also works that train and improve the retriever encoder (Zhang et al., 2023c), as well as distill and refine the data in the database (Jeronymo et al., 2023). Such retrieval frameworks have lower costs and can be more flexible in applications across different domains. Liu et al. (2023c) built upon a retrieval-enhanced framework and provided several plug-and-play modules, enabling its application in various professional fields.
2.6 Overall Development Status
@@
We select representative benchmarks to assess the current development of LLMs across education-related capabilities. Specifically, we mainly collect the results from three sources: Huggingface [^4] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, OpenCompass [^5] https://opencompass.org.cn/leaderboard-llm and C-Eval [^6] https://cevalbenchmark.com/static/leaderboard.html. The formal two are comprehensive leaderboards. C-Eval is a Chinese evaluation suite for foundation models spanning 52 diverse disciplines. We collect performance data from popular general LLMs on these benchmarks, and the compiled results are presented in Table 1, where one can observe that:
  • It is hard for a single LLM to be superior across all capabilities. GPT-4 exhibits the best overall performance among current LLMs. However, LLMs still lag significantly behind humans in some crucial abilities, such as achieving 94% accuracy in TruthfulQA (Lin et al., 2021), a performance level far superior to the optimal results of LLMs.
@@
Figure 3: Two frameworks towards LLM-based educational framework. (a) is the unified framework where one LLM model handles all the education-related questions. (b) is the MoE framework where the LLM controller is responsible for task distribution, and the corresponding expert model finishes the given task.
3 Potential of LLM-based Education System
LLMs can potentially transform online education by understanding a wide range of student questions, similar to human teachers. They aim to provide support across different subjects and skill levels. With the latest developments in LLMs, we suggest two approaches for creating LLM-based education systems. The first involves training a comprehensive LLM that can handle questions from various subjects. The second approach uses a mixture-of-experts (MoE) framework, integrating specialized models to support the system with an LLM controller to manage interactive dialogs with students.
3.1 Unified Approach
Research on whether LLMs can handle educational tasks has been underway. Wang and Demszky (2023) introduced three teacher coaching tasks for generative AI: (A) scoring transcript segments using classroom observation instruments, (B) identifying highlights and missed opportunities for effective instructional strategies, and (C) offering actionable suggestions to encourage more student reasoning. And evaluated by human teachers, ChatGPT on these tasks for elementary math classroom transcripts generates responses that are relevant to improving instruction, but they are often not novel or insightful. Beyond that, Phung et al. (2023) assessed the programming education ability of ChatGPT and GPT-4 by comparing them with human tutors. The result shows that GPT-4 performs way better than ChatGPT, even close to human tutors in some scenarios, while it also highlights some situations in which GPT-4 struggles. In particular, for the grading feedback and task creation scenarios that have a substantial gap in the performance of GPT-4 compared to that of human tutors.
3.2 MoE Approach
Section 2 reviewed the current development of LLMs across various capabilities. Unfortunately, despite the existence of comprehensive language models, such as GPT-4, these models often exhibit notable deficiencies in certain abilities. This situation poses a challenge, indicating that relying solely on an LLM itself for educational guidance involving all these capabilities is currently a difficult task. Yet, LLMs can achieve excellent results through fine-tuning individual capabilities, and their ability to comprehend human language is exceptionally strong. Therefore, we can aggregate models with distinct capabilities using a mixture-of-experts approach. By establishing an LLM-based controller for language interaction and task assignment with students, a currently feasible education system can be generated.
An education framework implemented with a mixture-of-experts (MoE) approach is illustrated in Figure 3(b), consisting of multiple models that excel in individual capabilities (not necessarily LLMs) and an LLM controller. The controller is mainly responsible for three tasks: (1) Understand the student’s request and decide which specific area or areas the request is about. (2) Re-form the request to fit the input of the specific areas’ expert models. (3) Aggregate the output of the related experts and generate the final response to the student.
The advantage of this approach is that training is less challenging, requiring training on individual capability-related datasets to obtain multiple models, each superior in a specific capability. However, the drawback lies in potential understanding errors during inference due to interactions. This includes errors introduced by the LLM controller when assigning tasks based on student input and integrating outputs from various experts. Currently, this approach might be a realistic way of achieving the LLM-based educational assistant system.
4 Future Directions
@@
Recently, more and more researchers have been trying to apply LLMs to handle education tasks, such as course design, student evaluation, lesson plan design, and others. Nevertheless, there are still many challenges and opportunities.
4.1 Interdisciplinary Reasoning Ability
Students may encounter interdisciplinary reasoning problems during real-world learning, requiring the education system to integrate multiple capabilities to formulate responses. As illustrated in Figure 1, the student intends to write a program to solve a mathematical problem, and the model needs first to comprehend the mathematical problem, devise a solution, and then generate the code. This process necessitates the model to synthesize both mathematical and programming capabilities.
@@ 4.1 Interdisciplinary Reasoning Ability
However, there is currently limited research in the integration of multiple interdisciplinary capabilities for LLMs at this stage, including both datasets and algorithms. Boyko et al. (2023) examined how LLMs augment scientific inquiry, code development, scientific writing process, etc., and they propose that LLMs can foster interdisciplinary work by bridging knowledge gaps across scientific fields. However, they mainly discuss the LLMs’ ability to help researchers’ interdisciplinary collaboration instead of their ability to answer interdisciplinary questions. Cultivating an LLM to obtain this ability would help to develop a unified education system, which is an essential research direction.
4.2 Student Modeling
Establishing an LLM-based education system allows students to articulate their personalized needs through dialogs. Through such conversations, we can extract or infer personalized features about students, such as their current mastery of topics and preferences in learning styles.
@@ 4.2 Student Modeling
Besides modeling students from conversations, some researches (Aher et al., 2023; Argyle et al., 2023) have shown that LLMs have certain abilities in simulating humans and generating human samples. Applying to education, this ability indicates a potential for LLM-based student simulation. In this way, for the students with few interaction records, the LLM-based simulator could generate more samples and provide data to help the expert model better understand the student. It could help human teachers develop teaching skills better.
4.3 Multi-modal Education
In education, multi-modal information is common, like geometry problems combining text and images or textbook concepts with illustrations. Building a general intelligent education system requires handling such multi-modal data. Thus, developing a multi-modal educational model is a promising future direction.
Notably, the development of multi-modal LLMs is rapidly advancing (Ye et al., 2023; Du et al., 2022). Different kinds of architectures and pre-train tasks are proposed (Du et al., 2022). However, the education domain often exhibits unique distribution characteristics in multi-modal information. Therefore, it may place higher demands on the model’s multi-modal reasoning capability.
5 Conclusion

### v3

@@ 3.6. Discussion
Table 1. Overview of LLMs’ performance on foundational education-related capabilities.
@@
GPT-4 OpenAI (2023b) 92.00 62.00 67.00 91.40 59.00 68.70
ChatGPT OpenAI (2023a) 57.10 48.60 48.10 79.50 47.00 54.40
TigerBot-70B-Chat-V2(70B) TigerResearch (2023) 54.36 61.30 30.50 82.83 75.40 -
LLaMA2(70B) Touvron et al. (2023b) 60.27 51.60 29.90 82.30 56.18 55.20
LLaMA(65B) Touvron et al. (2023a) 43.37 47.10 23.70 82.30 55.09 38.80
Yi(34B) lingyiwanwu (2023) 50.64 48.90 26.20 82.00 56.23 81.40
Vicuna(33B) TheVicunaTeam (2023) 13.72 44.90 15.20 83.00 56.16 39.80
WizardLM(30B) Xu et al. (2023) 34.42 - 26.08 76.32 49.14 -
Moss(16B) FudanUniversity (2023) 6.90 39.00 - 55.80 49.00 33.10
Qwen(14B) Bai et al. (2023) 58.98 52.7 43.90 80.20 49.43 72.10
Baichuan2(13B) Yang et al. (2023c) 55.30 51.50 17.07 66.90 48.98 40.00
Alpaca(7B) Taori et al. (2023) 0.15 39.50 9.10 75.71 36.28 29.90
ChatGLM3(6B) Zeng et al. (2022) 72.30 43.10 44.50 76.50 - 69.00
4. Overall Development Status
@@
Before exploring the possibilities of building an education system based on LLMs, we first need to investigate the performance of LLMs in capabilities related to education. We select representative benchmarks to assess the current development of LLMs across education-related capabilities. Specifically, we mainly collect the results from three sources: Huggingface [^3] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard, OpenCompass [^4] https://opencompass.org.cn/leaderboard-llm and C-Eval [^5] https://cevalbenchmark.com/static/leaderboard.html. The formal two are comprehensive leaderboards. C-Eval is a Chinese evaluation suite for foundation models spanning 52 diverse disciplines. We collect performance data from popular general LLMs on these benchmarks, and the compiled results are presented in Table 1, where one can observe that:
  • It is hard for a single LLM to be superior across all capabilities. Among the current LLMs, GPT-4 has shown the most impressive overall performance. However, utilizing GPT-4 comes with higher associated costs compared to other LLMs, which can be a significant consideration for users and organizations with limited budgets, and it has been surpassed by TigerBot in knowledge-based QA tasks. For mathematics, GPT-4 achieves optimal performance on the representative dataset GSM8K, but it exhibits a higher error rate in basic arithmetic tasks, such as large number multiplication.
  • LLMs still lag significantly behind humans in some crucial abilities. One notable example of this gap is illustrated by their performance on TruthfulQA (Lin et al., 2021), a benchmark designed to evaluate the ability of models to provide truthful and accurate answers, where human achieves achieving 94% accuracy while GPT-4 only got 59% correct.
@@
5. Potential of LLM-based Education System
LLMs can potentially transform online education by understanding a wide range of student questions, similar to human teachers. They aim to provide support across different subjects and skill levels. With the latest developments in LLMs, we suggest two approaches for creating LLM-based education systems. The first involves training a comprehensive and unified LLM that can handle questions from various subjects. The second approach uses a mixture-of-experts (MoE) framework, integrating specialized models to support the system with an LLM controller to manage interactive dialogues with students.
5.1. Unified Approach
Research on whether general LLMs can handle educational tasks has been underway. Wang and Demszky (2023) introduced three teacher coaching tasks for generative AI: (A) scoring transcript segments using classroom observation instruments, (B) identifying highlights and missed opportunities for effective instructional strategies, and (C) offering actionable suggestions to encourage more student reasoning. And evaluated by human teachers, ChatGPT on these tasks for elementary math classroom transcripts generates responses that are relevant to improving instruction, but they are often not novel or insightful. Beyond that, Phung et al. (2023) assessed the programming education ability of ChatGPT and GPT-4 by comparing them with human tutors. The result shows that GPT-4 performs way better than ChatGPT, even close to human tutors in some scenarios, while it also highlights some situations in which GPT-4 struggles. In particular, for the grading feedback and task creation scenarios that have a substantial gap in the performance of GPT-4 compared to that of human tutors.
From the methods proposed by researchers in developing LLMs for educational capabilities, we can extract some common, scalable approaches to lay the groundwork for developing a unified LLM-based educational system:
  • High-quality demonstrations. While collecting high-quality data from various fields for fine-tuning LLMs is impractical, achieving better responses through prompt engineering as a form of demonstration is feasible.
  • API Tool Learning. For inherent challenges within LLMs, such as large number computations and the absence of real-time information, these can be addressed by incorporating external APIs as tools.
  • Search-based methods. Attempts have been made across various fields to improve the task completion accuracy of LLMs using search-based methods, leveraging the probabilistic nature of LLMs. For challenging questions, LLMs might waver among multiple possible responses. Here, employing search-based methods to evaluate and filter all options can effectively enhance accuracy, offering a generally applicable solution.
The benefit of developing a unified LLM-based educational system is that it centers around a general LLM handling the core reasoning tasks. This setup means that all major language-based interactions are directly between the LLM and the students, making the system easier to deploy. The most significant effort and resources are invested during the training phase. This critical period is where the LLM gains expert-level skills in a wide range of subjects, preparing it to effectively support and educate students across various disciplines.
5.2. MoE Approach
Section  3 reviewed the current development of LLMs across various capabilities. Unfortunately, despite the existence of comprehensive language models, such as GPT-4, these models often exhibit notable deficiencies in certain abilities. This situation poses a challenge, indicating that relying solely on an LLM itself for educational guidance involving all these capabilities is currently a difficult task. Yet, LLMs can achieve excellent results through fine-tuning individual capabilities, and their ability to comprehend human language is exceptionally strong. Therefore, we can aggregate models with distinct capabilities using a mixture-of-experts approach. By establishing an LLM-based controller for language interaction and task assignment with students, a currently feasible education system can be generated.
An education framework implemented with a mixture-of-experts (MoE) approach is illustrated in Figure 4(b), consisting of multiple models that excel in individual capabilities (not necessarily LLMs) and an LLM controller. The controller is mainly responsible for three tasks:
  • Understand the student’s request and decide which specific area or areas the request is about.
  • Re-form the request to fit the input of the specific areas’ expert models.
  • Aggregate the output of the related experts and generate the final response to the student.
The advantage of the MoE approach is that training is less challenging. The result is a suite of models, each excelling in its specific domain or capability, which, when combined, offer a comprehensive educational tool. This specialization means training can be more focused and less onerous, optimizing resources towards developing excellence in distinct areas of knowledge and skills. However, one significant drawback is the increased potential for misunderstandings or errors during the system’s inference phase, primarily due to the complexity of interactions between the different specialized models and the LLM controller. Errors can arise from the LLM controller misinterpreting student inputs or incorrectly assigning tasks to the specialized models. Moreover, integrating outputs from various experts into a coherent response can also introduce discrepancies, as differences in context or terminology used by each model can lead to inconsistencies in the overall communication with students.
Despite these challenges, the approach represents a practical pathway toward realizing an LLM-based educational assistant system. By leveraging specialized models for different capabilities, it’s possible to create a more flexible and efficient system that can adapt to a wide range of educational needs and learning styles. The key to success lies in improving the integration and communication between the specialized models and the overarching LLM controller, ensuring that the system can handle complex inquiries and deliver accurate, useful responses to students. Currently, this approach appears to be a viable strategy for achieving the ambitious goal of an effective LLM-based educational assistant, promising a future where personalized education is accessible and adaptable to every learner’s need.
@@ 5.2. MoE Approach
Figure 4. Two frameworks towards LLM-based educational framework. (a) depicts the unified approach, where a single LLM addresses all aspects of educational-related queries, utilizing its internal capabilities such as mathematics, writing, knowledge-based question answering, reasoning, and programming. (b) illustrates the Mixture of Experts (MoE) approach, where an LLM controller is tasked with task distribution, delegating specific questions to specialized expert models that are proficient in individual areas.
6. Challenges and Future Directions
@@
Recently, more and more researchers have been trying to apply LLMs to handle education tasks, such as course design, student evaluation, lesson plan design, and others. Nevertheless, there are still numerous challenges and opportunities that need to be addressed.
  • Planning for Students. Solving subject-related questions for students can significantly address the issue of students not receiving targeted guidance from teachers. Furthermore, a higher-level task involves assessing students’ knowledge status and planning their learning paths. These tasks are continuously evolving in the era of deep learning, with the adaptation and application of LLMs in these areas requiring further exploration. The primary challenge in planning learning paths for students lies in integrating knowledge from two aspects: first, the human knowledge system, which involves the structural relationships between knowledge points, requiring LLMs to understand the meaning of these knowledge points. Second, the personalized information of students, including their knowledge state, learning interests, and habits. Previous deep learning models for this task have been trained on sequences of student behaviors, which are often constructed as IDs rather than text. Since the foundation of LLMs is their ability to process text, the gap in data form presents a significant challenge in applying LLMs to this task.
@@
  • Interdisciplinary Reasoning Ability. Students may encounter interdisciplinary reasoning problems during real-world learning, requiring the education system to integrate multiple capabilities to formulate responses. As illustrated in Figure 1, the student intends to write a program to solve a mathematical problem, and the model needs first to comprehend the mathematical problem, devise a solution, and then generate the code. This process necessitates the model to synthesize both mathematical and programming capabilities. However, there is currently limited research in the integration of multiple interdisciplinary capabilities for LLMs at this stage, including both datasets and algorithms. Boyko et al. (2023) examined how LLMs augment scientific inquiry, code development, scientific writing process, etc., and they propose that LLMs can foster interdisciplinary work by bridging knowledge gaps across scientific fields. However, they mainly discuss the LLMs’ ability to help researchers’ interdisciplinary collaboration instead of their ability to answer interdisciplinary questions. Cultivating an LLM to obtain this ability would help to develop a unified education system, which is an essential research direction.
@@
  • Student Modeling. Before the era of LLMs, in the age of deep learning, modeling student behavior was primarily achieved through sequential models, such as RNNs (Sherstinsky, 2020) and Transformers. A drawback of this approach was the inability to obtain student feedback, and the results lacked interpretability. Establishing an LLM-based education system allows students to articulate their personalized needs through dialog. Through such conversations, we can extract or infer personalized features about students, such as their current mastery of topics and preferences in learning styles. Besides modeling students from conversations, some researches (Aher et al., 2023; Argyle et al., 2023) have shown that LLMs have certain abilities in simulating humans and generating human samples. Applying to education, this ability indicates a potential for LLM-based student simulation. In this way, for the students with few interaction records, the LLM-based simulator could generate more samples and provide data to help the expert model better understand the student. It could help human teachers develop teaching skills better.
  • Social Bias of LLMs Even after training through Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), LLMs can to some extent avoid answers that do not align with human cultural habits and values (Jiang et al., 2022a; Feng et al., 2023b), it has been observed that LLMs still manifest a certain degree of value bias in their responses.  Feng et al. (2023b) pointed out that the training of LLMs can lead to a certain degree of political bias. In the field of education, although most questions posed by students are related to scientific knowledge, issues such as writing and text reasoning should ideally be avoided by researchers developing foundational LLM models. In educational applications, the social biases inherent in LLMs pose a risk of inadvertently impart imparting skewed value systems to students. To safeguard the educational integrity and ensure the neutral and fair dissemination of knowledge, it is crucial to implement stringent measures. These could include developing advanced content review systems, establishing clear guidelines for the ethical use of LLMs in educational settings, and continuously monitoring the quality and nature of the LLM-generated content. Through these efforts, the educational community can leverage the benefits of LLMs while minimizing the risk of perpetuating biases, thus maintaining a balanced and objective learning environment.
  • Preventing Cheating in Education. The texts generated by LLMs are indistinguishable from or even surpass those produced by humans in terms of fluency and usage. Although the primary aim of this article is to survey the development of LLMs in educational capacities, offering insights for the creation of an educational supermodel, it’s crucial to recognize that in certain educational contexts, the over-reliance on LLMs is not desirable as it could hinder the natural learning process. For instance, while it’s acceptable for students to seek assistance from LLMs to aid understanding during homework tasks, relying on LLMs to complete assignments without thoughtful consideration prevents students from receiving the necessary practice and learning. Therefore, identifying content generated by LLMs holds significant importance in the educational domain to prevent cheating and ensure the integrity of the learning process. By striking a balance between leveraging LLMs for educational enhancement and maintaining rigorous educational standards, educators and technologists can create an environment where students benefit from technology without compromising their learning journey. Recent studies have proposed detectors to identify LLM-generated texts. The foundational ideas could be broadly categorized into two primary strategies: statistical outlier detection methods and supervised classifiers. The former strategy focuses on uncovering statistical differences in linguistic features between texts written by humans and those generated by LLMs. This involves analyzing patterns, such as syntactic structures, vocabulary diversity, and stylistic nuances, that distinguish LLM-generated texts from human-crafted writings. These statistical indicators serve as markers for automated systems to detect content that deviates from human norms, potentially signaling LLM involvement. On the other hand, supervised classifiers rely on a different mechanism. This approach employs machine learning algorithms that have been trained on a labeled dataset containing examples of both human-written and LLM-generated texts. The battle against detecting LLM-generated texts is dynamic, necessitating ongoing research and adaptation of detection methodologies. As LLMs become increasingly sophisticated, the strategies for distinguishing their outputs from human-created content will need to evolve, embracing a combination of statistical insights, machine learning innovations, and perhaps new, yet-to-be-discovered methods.
  • Multi-modal Education. In education, multi-modal information is common, like geometry problems combining text and images or textbook concepts with illustrations. Building a general intelligent education system requires handling such multi-modal data. Notably, the development of multi-modal LLMs is rapidly advancing (Ye et al., 2023; Du et al., 2022). Different kinds of architectures and pre-train tasks are proposed (Du et al., 2022). However, the education domain often exhibits unique distribution characteristics in multi-modal information. Firstly, in education, images and text often have a high level of detail matching; for example, geometry questions often describe the specific parameters of shapes in images in great detail. Therefore, multi-modal large models need to have a high capability of capturing details in image information. Secondly, the multi-modal information in education often requires the model to have a high capacity for cross-modal reasoning, but such data is less common in multi-modal datasets, leading to a potential shortfall in the reasoning capabilities of multi-modal language models across different modalities. Addressing this gap may require targeted datasets, and inspired by Chain of Thought (CoT) and its variants, the data should ideally contain detailed steps of multi-modal reasoning. Efforts are currently being made to address these data deficiencies in the field. Moreover, the characteristics of image and text data in education could limit the choice of structures for multi-modal models. For instance, a popular approach in the field of multi-modal large models involves dividing images into patches to create “image tokens”, (McKinzie et al., 2024) which are then processed alongside text tokens as input. However, in the educational context, such division might disrupt certain key geometric structures within the images, thereby affecting their interpretation. This drawback could be more pronounced in educational multi-modal scenarios.
7. Conclusion